### v2

# Computationally-Efficient Linear Periodically Time-Variant Digital PLL Modeling Using Conversion Matrices and Uncorrelated Upsampling

Abstract
This paper introduces a conversion matrix method for linear periodically time-variant (LPTV) digital phase-locked loop (DPLL) phase noise modeling that offers precise and computationally efficient results to enable rapid design iteration and optimization. Unlike many previous studies, which either assume linear time-invariance (LTI) and therefore overlook phase noise aliasing effects, or solve LPTV systems with noise folding and multiple sampling rate conversions that heightens modeling and computational complexity, the proposed conversion matrix method allows the designer to represent the LPTV systems using intuitive LTI-like transfer functions with excellent accuracy. Additionally, computational efficiency is improved through the uncorrelated upsampling method, which eliminates the need to consider beat frequency of noise sources with different sampling rates. The proposed algorithm is applied to modeling a DPLL with time-varying proportional loop gain, and the modeling accuracy is validated with Simulink transient simulations.
Index Terms:
Conversion matrix, harmonic transfer matrix (HTM), phase-locked loop (PLL), phase noise, linear periodically time-variant (LPTV), noise folding.
I Introduction
Contemporary System-on-Chip (SoC) designs typically require the integration of multiple frequency synthesizers to deliver precise clocks for mixers and digital circuits. Such circuits can end up occupying large physical area (e.g., 25% of the total chip area in [1]), and can end up limiting the performance of such systems. Digital phase-locked loops (DPLLs) are a compelling option for frequency synthesis, due to their technology-scalable area, high programmability, and excellent noise-power figure-of-merit (FoM) [2, 3, 4, 5, 6, 7, 8, 9].
Designing high-performance DPLLs necessitates a comprehensive understanding of the output phase noise (PN). To determine the transfer function from each noise source to the output, many prior studies have treated PLLs as linear time-invariant (LTI) systems at the reference frequency, $f_{ref}$, neglecting noise aliasing caused by divider downsampling [10, 11, 12, 13, 14, 15, 16]. This LTI approximation is suitable for PLLs with relatively low bandwidth. However, for PLLs with very wide bandwidth to suppress the excessive PN from ring oscillators, e.g., 80 MHz in [17] and $\sim 100$ MHz in [18], the aliased noise will contribute significantly to the in-band PN and thus cannot be ignored. Additionally, [12] uses forward Euler mapping $z=1+sT_{ref}$ for $z$-domain to $s$-domain conversion, making the model less accurate in predicting PN close to $f_{ref}/2$.
For a precise modeling of PN in wide-bandwidth DPLLs and multiplying delay-locked loops (MDLLs), it’s essential to consider the system as a discrete-time linear periodically time-variant (LPTV) system. The methodologies in [19, 20, 21] provides great accuracy on their respective models under analysis. However, these math-intensive, topology-specific derivations hinder engineers from quickly adapting to increasingly complicated new topologies [20, 21]. The complexity increases further when a delta-sigma modulator (DSM) with a distinct clock frequency from the oscillator and reference is incorporated [19], necessitating multiple sample-rate conversions within the loop.
The conversion matrix method, also known as the harmonic transfer matrix (HTM) method [22], offers an alternative approach for solving transfer functions in LPTV systems. This approach, previously employed in mixer/$N$-path filter analysis [23, 24, 25], has demonstrated simplified yet accurate results compared to traditional state-space equation solutions [26, 27, 28, 29]. Moreover, it has been leveraged in stability analysis of grid-connected PLLs in power converters [30, 31, 32]. The primary advantage offered by the conversion matrix method is its capability to represent all LPTV components with matrices, where the elements denote the ratio of the input and output frequency-shifted spectrum. Consequently, formulating the problem is straightforward for designers, and computationally demanding matrix operations are offloaded to modern computers optimized for such calculations. The conversion matrix method provides a modeling framework that can be easily applied to different designs with arbitrary complexity.
PLLs, which are inherently LPTV, appear to be an ideal candidate for the conversion matrix method. However, the incorporation of a DSM could significantly increase the computational complexity. To understand why, consider simulating the conversion gain of an $RC$-mixer using Spectre periodic steady-state (PSS) and periodic AC (PAC) analysis. When the clock frequency, $f_{s}$, is 1 GHz, the fundamental frequency in PSS equals to $f_{s}=1$ GHz; the simulation can quickly converge and only a few sidebands, depending on the application, need to be saved. However, incorporating an additional mixer with a clock frequency $f_{s}^{{}^{\prime}}=f_{s}+\Delta f=$ 1.01 GHz changes the system’s fundamental frequency to the beat frequency $\Delta f=$ 0.01 GHz. As a result, 100$\times$ more sidebands need to be saved, significantly increasing the simulation memory and time consumption. Similar challenges arise when using the conversion matrix method to model a PLL if the DSM frequency is not an integer multiple of the reference frequency. Here, the beat frequency becomes the fundamental frequency, potentially making the problem dimension of the conversion matrix method (similar to the number of sidebands in PSS) too large to be solved.
The contributions of this paper are as follows:
  1. 1.
    A ground-up theory for discrete-time LPTV system noise analysis using the conversion matrix method, accommodating multiple different sampling rates for independent noise sources.
  2. 2.
    Introducing uncorrelated upsampling that allows ignoring any beat frequency between different-rate noise sources, thereby reducing computation time.
  3. 3.
    Development of a modularized algorithm enabling easy adaptation to different PLL designs. The MATLAB code and simulation model can be downloaded from [33].
The paper is organized as follows: Section II defines the conversion matrix method in discrete-time LPTV systems. Section III introduces uncorrelated upsampling and provides an example of modeling a PLL with time-varying proportional loop gain. Section IV compares the simulation results with the PN predicted by the model. Finally, Section V offers conclusions.
II Discrete-Time LPTV System Modeling
In a discrete-time LPTV system depicted in Fig. 1, the relationship between input $x[n]$ and output $y[n]$ is characterized by the time-varying impulse responses of the feed-forward and feedback paths, represented as $p_{n}[n]$ and $f_{n}[n]$, respectively. The LPTV system’s frequency conversion behavior, with a sampling rate of $f_{s}=1/T_{s}$, can be eventually expressed first by using a vector, $\mathbf{X}$, composed of the frequency-shifted input spectrum:
$\mathbf{X}=\begin{bmatrix}X(\Omega-(K-1)\frac{2\pi}{K})\\ \vdots\\ X(\Omega-k\frac{2\pi}{K})\\ \vdots\\ X(\Omega)\end{bmatrix},$ (1)
in which $X(\Omega)$ is the discrete-time Fourier transform (DTFT) of $x[n]$ [34, eq. (12.1)], $\Omega=2\pi fT_{s}$ is the discrete angular frequency, $K$ is defined as the problem dimension, $k$ is the harmonic index of the input, and $2\pi/K$ is the fundamental discrete angular frequency of the time-varying impulse response.
Similarly, define the output vector, $\mathbf{Y}$, by replacing $X$ with $Y$ in (1),
$\mathbf{Y}=\begin{bmatrix}Y(\Omega-(K-1)\frac{2\pi}{K})\\ \vdots\\ Y(\Omega-q\frac{2\pi}{K})\\ \vdots\\ Y(\Omega)\end{bmatrix},$ (2)
in which $q$ is the harmonic index of the output. The output $\mathbf{Y}$ can be written as a matrix, $\mathbf{H}$, multiplied by the input $\mathbf{X}$:
$\mathbf{Y}=\mathbf{H}\mathbf{X},$ (3)
in which
$\mathbf{H}=\begin{bmatrix}H_{K-1,K-1}&\ldots&H_{K-1,k}&\ldots&H_{K-1,0}\\ \vdots&\ddots&\vdots&&\vdots\\ H_{q,K-1}&\ldots&H_{q,k}&\ldots&H_{q,0}\\ \vdots&&\vdots&\ddots&\vdots\\ H_{0,K-1}&\ldots&H_{0,k}&\ldots&H_{0,0}\\ \end{bmatrix}$ (4)
is defined as conversion matrix that relates frequency-shifted input and output spectrum. The elements of $\mathbf{H}$, denoted as $H_{q,k}$, represent the gain from the $k^{\text{th}}$ input harmonic to the $q^{\text{th}}$ output harmonic. The discrete-time conversion matrix has a finite $K\times K$ dimension, as opposed to being a doubly infinite matrix in the continuous-time case [22, eq. (12)]. This is because the DTFT of a discrete-time signal is $2\pi$-periodic, making the definitions in (1)-(4) contain all possible combinations of the input and output harmonics.
If the input $x[n]$ is a wide-sense stationary (WSS) stochastic process, the double-sideband power spectral density (PSD) of the output $y[n]$ can be written as [35]:
$S_{yy}(\Omega)=\sum_{k=0}^{K-1}|H_{0,k}(\Omega)|^{2}Sxx\left(\Omega-\frac{2\pi k}{K}\right).$ (5)
The target now is to determine $\mathbf{H}$ for the system in Fig. 1, so (5) can be used to calculate the output PSD. By inspection,
$\mathbf{Y}=\mathbf{P}(\mathbf{X}-\mathbf{F}\mathbf{Y}),$ (6)
in which $\mathbf{P}$ and $\mathbf{F}$ are conversion matrices associated with $p_{n}[n]$ and $f_{n}[n]$, respectively. Rearranging the terms in (6), the output $\mathbf{Y}$ is
$\mathbf{Y}=[(\mathbf{I}+\mathbf{P}\mathbf{F})^{-1}\mathbf{P}]\mathbf{X}=\underbrace{[(\mathbf{I}+\mathbf{P}\mathbf{F})\backslash\mathbf{P}]}_{\mbox{$\equiv\mathbf{H}$}}\mathbf{X},$ (7)
[…]
Within an LPTV system, the input $x[n]$ can go through 2 types of building blocks: LTI subsystems, and multiplication with a periodic sequence. Subsequent sections will derive the conversion matrices for both blocks, ultimately leading to the calculation of $\mathbf{H}$.
Figure 1: Linear periodically time-variant (LPTV) system with input $x[n]$ and output $y[n]$.
II-A LTI Subsystem
The discrete-time LTI system depicted in Fig. 2 is characterized by a time-invariant impulse response, $h[n]$, and its corresponding DTFT, $H(\Omega)$. The relation between the input and output is given by:
$Y(\Omega)=H(z)\bigg{|}_{z=e^{j\Omega}}X(\Omega),$ (8)
in which $H(z)$ is the $z$-transform of $h[n]$. By extending the scalar operation in (8) to the matrix operation described in (3), the output $\mathbf{Y}$ can be represented as:
$\mathbf{Y}=\underbrace{\begin{bmatrix}H(\Omega-\frac{2\pi(K-1)}{K})&&&&\\ &\ddots&&&\\ &&H(\Omega-\frac{2\pi k}{K})&&\\ &&&\ddots&\\ &&&&H(\Omega)\\ \end{bmatrix}}_{\mbox{$\equiv\mathbf{H}_{\text{LTI}}$}}\mathbf{X}.$ (9)
Here, $\mathbf{H}_{\text{LTI}}$ is the conversion matrix of the LTI system associated with the impulse response $h[n]$. The absence of frequency conversion in an LTI system renders $\mathbf{H}_{\text{LTI}}$ a $K\times K$ diagonal matrix.
Figure 2: Linear time-invariant (LTI) system with input $x[n]$ and output $y[n]$.
II-B Multiplication with Periodic Sequence
Multiplying the input, $x[n]$, with a periodic sequence, $w[n]$, shown in Fig. 3 is another fundamental block in LPTV systems. The sequence $w[n]$ has a period of $N$, i.e., $w[n]=w[n+N]$. Define $w_{0}[n]$ as one period of $w[n]$ [34, eq. (12.17)]
$w_{0}[n]=\begin{cases}w[n],&0\leq n\leq N-1\\ {0},&{\text{else}}\end{cases}.$ (10)
The DTFT of $w[n]$ can be written as [34, eq. (12.23)]
$W(\Omega)=\frac{2\pi}{N}\sum_{k=-\infty}^{\infty}W_{0}\left(\frac{2\pi k}{N}\right)\delta\left(\Omega-\frac{2\pi k}{N}\right),$ (11)
in which $W_{0}(\Omega)$ is the DTFT of $w_{0}[n]$ [34, eq. (12.1)]
$W_{0}(\Omega)=\sum_{n=0}^{N-1}w_{0}[n]e^{-jn\Omega}.$ (12)
Multiplication in the time domain translates to circular convolution in frequency domain [34, eq. (12.14)]. As a result, the output spectrum $Y(\Omega)$ can be written as
$Y(\Omega)$ $=\textbf{DTFT}\{x[n]w[n]\}$ (13)
$=\frac{1}{N}\sum_{k=-\infty}^{\infty}W_{0}\left(\frac{2\pi k}{N}\right)X\left(\Omega-\frac{2\pi k}{N}\right).$
Using (13), the output $\mathbf{Y}$ becomes
$\mathbf{Y}=\underbrace{\begin{bmatrix}H_{0}&\ldots&H_{1-j}&\ldots&H_{1-N}\\ \vdots&\ddots&\vdots&&\vdots\\ H_{i-1}&\ldots&H_{0}&\ldots&H_{i-N}\\ \vdots&&\vdots&\ddots&\vdots\\ H_{N-1}&\ldots&H_{N-j}&\ldots&H_{0}\\ \end{bmatrix}}_{\mbox{$\equiv\mathbf{H_{\text{Mul}}}$}}\mathbf{X},$ (14)
where $\mathbf{H}_{\text{Mul}}$ is the conversion matrix corresponding to the multiplication of the periodic sequence with the input. The element of $\mathbf{H}_{\text{Mul}}$ in the $i^{\text{th}}$ row and $j^{\text{th}}$ column is given by
$H_{\text{Mul}(i,j)}=\frac{1}{N}W_{0}\left(\frac{2\pi(i-j)}{N}\right).$ (15)
Like its continuous-time counterpart, $\mathbf{H}_{\text{Mul}}$ is a $K\times K$ frequency-independent Toeplitz matrix [22, eq. (27)].
Figure 3: Multiply the input $x[n]$ by a periodic sequence $w[n]$ and generate the output $y[n]$.
III PLL Modeling Example
With the derived building blocks, this section models a digital PLL with fast phase-error correction (FPEC) [7] as an example. In contrast to a conventional PLL that maintains a constant proportional loop gain $K_{P}$, FPEC boosts $K_{P}$ at the beginning of each reference cycle to rapidly eliminate the residue jitter from the previous reference cycle [7, Fig. 2], resulting in reduced average jitter. This is a time-varying concept, and cannot be modeled with conventional LTI techniques.
The high-level block diagram of the Type-II DPLL with FPEC is shown in Fig. 4. The $N$-periodic time varying loop gain $K_{P}[n]$ can be written as:
$K_{P}[n]=\begin{cases}\frac{NK_{P0}}{P},&kN\leq n\leq kN+P-1\\ {0},&{\text{else}}\end{cases},$ (16)
in which $k\in\{0,1,2,\ldots\}$, $P\in[1,N]$ is the length of the FPEC window, and $K_{P0}$ is the proportional loop gain when FPEC is deactivated, i.e., $P=N$. When $P$ decreases, the height of $K_{P}[n]$ increases while keeping the total area within one reference period the same. Subsequent subsections will derive the DCO-rate transfer functions for noise sources with different sampling rates.
Figure 4: High-level block diagram of the PLL with time-varying proportional loop gain $K_{P}[n]$.
III-A Noise Sampling Rate Conversion
All noise sources must be converted to the DCO rate before being fed into the DCO-rate PLL model. As depicted in Fig. 5, the noise sources operate at 3 different rates:
  1. 1.
    DCO noise, $\phi_{n,DCO}[n]$, running at $f_{DCO}$, which is the highest rate in the model.
  2. 2.
    DSM quantization noise, $q_{n,DSM}[p]$, running at $f_{DSM}=f_{DCO}/M$. In most cases $1\leq M<N$ for pushing the shaped quantization noise to higher frequency.
  3. 3.
    Reference noise, $\phi_{n,ref}[m]$, running at $f_{ref}=f_{DCO}/N$, which ideally should be the fundamental frequency of the model, making $K=N$ in (1). The phase detector (PD) quantization noise can be lumped into the reference noise by dividing PD gain $K_{PD}$.
Among these sources, DCO noise can be directly fed into the DCO-rate PLL model, while the sampling rate of the reference and DSM noise necessitates upconversion to the DCO rate.
The sample-rate upconversion is achieved through uncorrelated upsampling. Uncorrelated upsampling by a factor of $L$ involves upsampling by inserting $L-1$ zeros between adjacent samples of the input $x[m]$, followed by multiplying a decorrelation sequence, $w_{U,L}[n]$, that removes correlated frequency spectrum with respect to the fundamental discrete angular frequency, $2\pi/N$. The PSD of the upsampled sequence $y_{corr}[n]$, $S_{yy,corr}(\Omega)$, can be written as
$S_{yy,corr}(\Omega)=\frac{1}{L}S_{xx}\biggl{(}\text{mod}(L\Omega,2\pi)\biggr{)},\Omega\in[0,2\pi),$ (17)
in which $\text{mod}(a,b)$ finds the remainder after division of $a$ by $b$. One period of $w_{U,L}[n]$ is expressed as
$w_{U0,L}[n]=\begin{cases}\sqrt{\mathrm{gcd}(L,N)},&\text{if mod}(n,\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases},$ (18)
in which $10\leq n\leq N-$ 1, while $\mathrm{gcd}(L,N)$ computes the greatest common divisor of $L$ and $N$. The decorrelation sequence $w_{U,L}[n]$ is the $N$-periodic extension of $w_{U0,L}[n]$:
$w_{U,L}[n]=w_{U0,L}[\text{mod}(n,N)],$ (19)
and its associated conversion matrix $\mathbf{W}_{U,L}$ can be found with (15). An intuitive explanation of $w_{U,L}[n]$ is shown in the Appendix.
Note that multiplying the upsampled noise source with a decorrelation sequence $w_{U,L}[n]$ does not occur in an actual PLL; it only has mathematical significance that allows using $f_{ref}$ as the fundamental frequency of the system. Otherwise, any beat frequency lower than $f_{ref}$ becomes the fundamental frequency, unnecessarily enlarging the problem dimension as $K$ increases in (1). For example, if $N=51$ and $M=50$, the problem dimension $K=\text{lcm}(M,N)=2550$, where lcm represents the least common multiple. Considering the $O(n^{3})$ complexity when inverting an $n\times n$ matrix, the proposed uncorrelated upsampling reduces the computation time by $(\text{lcm}(M,N)/N)^{3}=125000\times$.
Figure 5: Multi-rate LPTV PLL model for time-varying loop gain analysis. All noise sources are converted to DCO rate with uncorrelated upsampling for further processing.
III-B DCO-Rate Transfer Function Analysis
Figure 6: DCO-rate LPTV model for FPEC PLL.
After converting all noise sources to the DCO rate, noise transfer functions are derived utilizing Fig. 6. The divider is modeled as multiplying the output phase by a $N$-periodic pulse, $w_{R,1}[n]/N$, that samples the output phase every $N$ samples:
$w_{R,1}[n]=\begin{cases}1,&n=kN\\ {0},&{\text{else}}\end{cases}$ (20)
for $k\in\{0,1,2,\ldots\}$. The time-varying loop gain implements (16) by multiplying a constant $NK_{P0}/P$ with a $N$-periodic window $w_{R,P}[n]$:
$w_{R,P}[n]=\begin{cases}1,&kN\leq n\leq kN+P-1\\ {0},&{\text{else.}}\end{cases}$ (21)
for $k\in\{0,1,2,\ldots\}$. Multiplying with (20) or (21) falls under the case depicted in Fig. 3, and the associated conversion matrices,$\mathbf{W}_{R,1}$ and $\mathbf{W}_{R,P}$, can be calculated using (15).
The conversion matrices for the following LTI transfer functions can be obtained using (9):
  1. 1.
    $\mathbf{H}_{ZOH,N}$: conversion matrix of $N$-length zero-order hold $H_{ZOH,N}(z)=\frac{1-z^{-N}}{1-z^{-1}}$.
  2. 2.
    $\mathbf{K}_{I}$: conversion matrix of the integral path transfer function $\frac{K_{I}}{1-z^{-1}}$.
  3. 3.
    $\mathbf{H}_{DCO}$: conversion matrix of the DCO transfer function $H_{DCO}=\frac{K_{DCO}}{f_{DCO}}\frac{z^{-1}}{1-z^{-1}}$. $K_{DCO}$ is the DCO gain with unit Hz/LSB.
The conversion matrices for $\phi_{n,ref,up}[n]$, $q_{n,DSM,up}[n]$, and $\phi_{n,DCO}[n]$ to the output $\phi_{n,PLL}[n]$ are
$\begin{split}\mathbf{H}_{ref,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{ref}\\ \mathbf{H}_{DSM,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{DSM}\\ \mathbf{H}_{DCO,tot}&=(\mathbf{I}-\mathbf{L})\backslash\mathbf{P}_{DCO}\end{split}$ (22)
respectively, in which $\mathbf{L}$ is the LTI-like loop gain
$\mathbf{L}=-K_{PD}\mathbf{H}_{DCO}\underbrace{\left(\frac{NK_{P0}}{P}\overbrace{\mathbf{W}_{R,P}}^{\text{FPEC}}\mathbf{H}_{ZOH,N}+\mathbf{K}_{I}\right)}_{\text{Loop Filter}}\overbrace{\frac{\mathbf{W}_{R,1}}{N}}^{\text{Divider}}$ (23)
and
$\begin{split}\mathbf{P}_{ref}&=K_{PD}\mathbf{H}_{DCO}\left(\frac{NK_{P0}}{P}\mathbf{W}_{R,P}\mathbf{H}_{ZOH,N}+\mathbf{K}_{I}\right)\\ \mathbf{P}_{DSM}&=\mathbf{H}_{DCO}\mathbf{H}_{ZOH,N}\\ \mathbf{P}_{DCO}&=\mathbf{I}\end{split}$ (24)
are path gains from each input noise source to the output. Together with (5) and uncorrelated upsampling shown in Fig. 5, the output phase noise, $\mathcal{L}(\Omega)$, can be determined as [36]:
$\mathcal{L}(\Omega)=\frac{1}{2}S_{\phi,PLL}(\Omega),$ (25)
in which $S_{\phi,PLL}(\Omega)$ is the SSB PSD of $\phi_{n,PLL}[n]$.
The process of finding the PLL output phase noise is summarized in Algorithm 1.
Algorithm 1 Finding Output PN Using Conversion Matrix.
Input: $S_{xx}(\Omega)$, $L$
Output: $\mathcal{L}(\Omega)$
$S_{xx,corr}(\Omega)\leftarrow S_{xx}(\Omega)\uparrow L$triangleright Upsampling by $L$, (17)
Compute $\mathbf{W}_{R,1}$, $\mathbf{W}_{R,P}$, $\mathbf{W}_{U,L}$triangleright Window CM, (15)
for $i=1:\text{length}(\mathbf{\Omega}_{vec})$ do
     $\Omega_{0}\leftarrow\mathbf{\Omega}_{vec}(i)$
     Evaluate $\mathbf{H}_{LTI}(\Omega_{0})$triangleright (9)
     $\mathbf{L}(\Omega_{0})\leftarrow f(\mathbf{W}_{R},\mathbf{H}_{LTI}(\Omega_{0}))$triangleright Loop gain, (23)
     $\mathbf{P}(\Omega_{0})\leftarrow g(\mathbf{W}_{R},\mathbf{H}_{LTI}(\Omega_{0}))$triangleright Path gain, (24)
     $\mathbf{H}(\Omega_{0})\leftarrow(\mathbf{I}-\mathbf{L}(\Omega_{0}))\backslash\mathbf{P}(\Omega_{0})$
     $\mathbf{H}_{U,L}(\Omega_{0})\leftarrow\mathbf{H}(\Omega_{0})\mathbf{W}_{U,L}$triangleright Decorrelation, (19)
     $\mathcal{L}(\Omega_{0})\leftarrow h(\mathbf{H}_{U,L}(\Omega_{0}),S_{xx,corr}(\Omega_{0}))$triangleright (5)
end for
In Algorithm 1, $\mathbf{\Omega}_{vec}$ contains frequency values of interest sampled within $[0,\pi)$, and the algorithm solves a $N\times N$ linear system of equations each time with all LTI transfer functions evaluated at $\Omega_{0}=\mathbf{\Omega}_{vec}(i)$. Compared with simulation-based phase noise extraction, the proposed algorithm computes $\mathcal{L}(\Omega_{0})$ almost instantaneously, delivering accuracy comparable to transient simulations that might be time-consuming.
IV Simulation Results
The proposed algorithm, which calculates the output PN using conversion matrices, is validated with Simulink transient simulations. The Simulink model implements Fig. 6 with parameters shown in Table I. The TDC resolution, $\Delta t_{res}$, is set to 15.16 ps, leading to $K_{PD}=1/(2\pi f_{ref}\Delta t_{res})$ = 300 $\text{rad}^{-1}$ referring to the reference frequency $f_{ref}=35$ MHz. The reference is assumed to have much less phase noise than the TDC, and is thus neglected in the following simulations.
TABLE I: Simulink Simulation Parameters
Parameter Value
N 18
$K_{PD}$ 300 $\text{rad}^{-1}$
$K_{P0}$ 0.4
$K_{I}$ $K_{P0}/32$
$K_{DCO}$ 4 MHz/LSB
$f_{ref}$ 35 MHz
IV-A DCO
Fig. 7 illustrates results when only the DCO noise source is active. The DCO noise is modeled as a white noise passing through a discrete-time integrator. The calculated output PN matches the simulated PN extremely well for both $P=N=18$ and $P=2$, including dips located at integer multiples of the reference frequency. The FPEC can suppress the in-band PN by 3.3 dB, while the PN above 20 MHz is increased.
The effect of FPEC window length $P$ on in-band PN and total integrated jitter is shown in Fig. 8. When $P$ reduces from 18 to 1, the in-band phase noise reduces 3.6 dB with a slope of 0.21 dB, and the RMS jitter reduces from 20.7 ps to 16.2 ps with a slope of 0.26 ps.
With excellent modeling accuracy provided above, the elapsed time of the proposed algorithm is 0.67 s for 16000 frequency domain samples, 21$\times$ shorter than the Simulink simulation time consumption. The efficiency of the algorithm can be further leveraged if only a few, instead of 16000, spot PN needs to be optimized, allowing the designer to quickly explore the design space.
Figure 7: Simulated output PN due to DCO noise with FPEC window length $P=2$ and without FPEC ($P=18$).
Figure 8: PN and jitter improvement with different FPEC window length $P$.
IV-B DSM
The PLL output PN due to DSM quantization noise is shown in Fig. 9. The DSM quantization noise is modeled as a white noise with power of $1/12$ passing through $(1-z^{-1})^{2}$. For both $M=3$ and $M=16$ cases, the proposed model, shown in the solid black lines, precisely predict the simulated PN shown in the blue and green lines. The decorrelation sequence (19) corrects the 3 dB lower in-band PN shown in the dashed black lines where the decorrelation sequence is not used. The high-frequency PN shows approximately 1 dB improvement in accuracy as well.
Figure 9: Simulated output PN due to DSM quantization noise for DSM clock divide ratio $M=3$ and $M=16$. The FPEC window length $P$ is the same as reference clock divide ratio $N$ = 18.
Contrary to the continuous-time analysis predicting that the DSM noise has a high-pass shape (similar to the dashed green and blue lines) in the output PN, the actual low-frequency PN is dominated by aliased DSM noise. The output PN at 1 MHz and the integrated jitter is shown in Fig. 10 with different DSM clock divide ratio, $M$, for $N$ = 18 and $N$ = 16. With smaller $M$, i.e., larger $f_{DSM}$, the quantization noise is pushed to higher frequency, resulting in a general trend of smaller RMS jitter. The 1 MHz spot PN, however, first increases with larger $f_{DSM}$, peaks at $M/2$, then decreases. When the in-band PN is critical in the application, $M$ needs to be carefully selected to avoid excessive noise aliasing. If the above results were to be generated without uncorrelated upsampling, the time consumption will be much larger under certain cases, e.g., 125$\times$ for $N=18$ and $M=5$.
Certain $M$ values lead to higher aliased noise, as seen in the peaks of Fig. 10. To find these numbers, we can factor $N$ into multiplication of prime numbers. For example, $N=18=2\times 3\times 3$. If $M$ can be factored to prime numbers within the same set $\{2,3\}$, then this $M$ will generate higher PN and jitter due to aliasing. For example, $M=2,3,4,6,9,10,12,14,15,16,18$. Among these numbers, $M=3,6,9,12,15$ generates an even higher noise floor, as 3 occurs twice when factoring $N$. Similarly, for $N=16=2\times 2\times 2$, any even $M$ generates more aliased PN. These $M$ values are preferred to be avoided when choosing the DSM operating frequency given a reference divide ratio $N$.
Figure 10: Simulated output PN and output RMS jitter due to DSM quantization noise with different DSM clock divide ratio $M$ for (a) $N$ = 18 and (b) $N$ = 16.
IV-C TDC
The simulated output PN due to TDC quantization noise is shown in Fig. 11 for FPEC window length $P=2$ and in the absence of FPEC ($P=18$). The TDC input-referred quantization noise floor is $\mathcal{L}_{TDC,in}=1/(12)/K_{PD}^{2}/f_{ref}=-135.8$ dBc/Hz. The PLL output PN has a low-pass shape with a low-frequency plateau $25.120\log N=$ 25.1 dB higher than the TDC input-referred noise floor. The colored simulation results matches well with the black lines predicted by the model. In the case of $P=2$, the output jitter due to TDC is 3.43 ps, 1 dB higher than 3.04 ps when $P=18$.
Figure 11: Simulated output PN due to TDC quantization noise for FPEC window length $P=2$ and without FPEC ($P=N=18$).
Fig. 12 shows the output RMS jitter due to TDC quantization noise for different FPEC window length $P$. Using FPEC with shorter pulse width $P$ increases the reference noise contribution while suppressing more DCO noise, as shown in Fig. 7. This trade-off is leveraged by using an optimal-threshold TDC (OT-TDC) [7] that generates less quantization noise by providing 2 more quantization thresholds than the bang-bang PD with optimal spacing.
Figure 12: Output RMS jitter due to TDC quantization noise for different FPEC window length $P$.
IV-D Total Output PN
Fig. 13 shows the output PN with all noise sources activated: TDC quantization noise, DSM quantization noise, and DCO phase noise. When $P$ reduces from 18 to 2, the in-band PN shows a 2.2 dB improvement, while the out-of-band noise plateau due to DSM quantization noise rises by 1.3 dB. Fig. 14 shows the in-band PN and RMS jitter improvement in finer resolution of $P$. The integrated RMS jitter improves only 0.93 dB when $P$ reduces from 18 to 2, a result of the trade-off between lower DCO noise and higher reference/TDC noise.
Figure 13: Total output PN with TDC quantization noise, DSM quantization noise, and DCO phase noise for $N=18$, $M=4$, and $P=18,2$.
Figure 14: Output PN and RMS jitter with TDC quantization noise, DSM quantization noise, and DCO phase noise for different $P$ under $N=18$ and $M=4$.
IV-E Algorithm Complexity
Fig. 15 shows the computation time required for a single frequency-domain sample using the proposed conversion matrix-based algorithm across varying problem dimensions $N$, specifically when $M=N-1$ and $P=2$. This algorithm was evaluated on a laptop equipped with an AMD Ryzen 9 5900HS CPU and 16.0 GB of RAM. For values where $N<150$, computation time exhibits a relationship proportional to $N^{2}$. However, for greater dimensions ($N\geq 200$), the complexity appears to be proportional to $N^{2.5}$. The proposed uncorrelated upsampling approach can reduce computation time by a factor of at least $(N-1)^{2}$. When $N<65$, the computation time for one frequency-domain sample remains under 1 ms, enabling sweeps across the design space to fine-tune PN at selected frequencies with great accuracy.
Figure 15: Computation time for one frequency-domain sample with different problem dimensions. The green dots are actual computation time and the fitted curves are shown in black.
V Conclusion
This paper introduces the conversion matrix method into modeling a discrete-time LPTV system, such as DPLLs. The conversion matrices for two fundamental building blocks, namely the LTI subsystem and multiplication with periodic sequence, are derived. In the DCO-rate model, all noise sources are uncorrelated upsampled to the DCO frequency, with the decorrelation sequence allowing the omission of the beat frequency of different noise sources. As a result, the computation resource consumption is significantly reduced. The proposed modeling method is verified against Simulink transient simulations, demonstrating high accuracy in predicting the phase noise.
[Intuitive Explanation of $w_{U,L}$]
Figure 16: Illustration of spectrum with sampling rate conversion. (a) The input spectrum at low sampling rate, (b) the spectrum after upsampling, and (c) the frequency-shifted spectrum after upsampling.
This Appendix provides the intuition that helps finding the sequence $w_{U,L}[n]$. In Fig. 5, $x[m]$ is a WSS low sampling rate sequence that needs to be upsampled. The DTFT of $x[m]$, $X(\Omega)$, is shown in Fig. 16(a). The height inside the shaded area at a specific $\Omega$ contains information on correlation; having different shaded height means uncorrelated spectrum. For example, $X(a)$ is uncorrelated with $X(b)$, i.e., $E\{X(a)X(b)\}=0$, because they have different shaded height. On the other hand, $X(a)$ is correlated with $X(c)$, i.e., $E\{X(a)X(c)\}\neq 0$, for having the same shaded height. When $x[m]$ is a real WSS random process, $X(\Omega)$ is conjugate symmetric w.r.t. $\Omega=k\pi$ and aperiodic for $\Omega\in[0,2\pi)$.
After upsampling by $L$ times by inserting $L-1$ zeros between adjacent samples, $y_{corr}[n]$ is no longer a WSS random process, but a cyclostationary random process. The autocorrelation sequence of $y_{corr}[n]$, $R_{yy,corr}$, is $L$-periodic, making the period of $Y_{corr}(\Omega)$ equal to $2\pi/L$. Fig. 16 shows $Y_{corr}(\Omega)$ when $L=6$.
When $y_{corr}[n]$ passes through a system with fundamental frequency $2\pi/N$, the output spectrum of the system $Y_{sys}(\Omega)$ is a summation of frequency-shifted $Y_{corr}(\Omega)$:
$Y_{sys}(\Omega)=\sum_{k=0}^{N-1}H_{0,k}(\Omega)Y_{corr}\left(\Omega-\frac{2\pi k}{N}\right).$ (26)
Fig. 16(c) shows $Y_{corr}(\Omega-\pi/4)$ for $N=8$ and $k=1$. By inspection, $Y_{corr}(\Omega-\pi/4)$ is uncorrelated with $Y_{corr}(\Omega)$. In general, there are 2 types of frequency shifts:
  1. 1.
    Correlated shift: the shifted spectrum is correlated with the original spectrum. For $L=6$ and $N=8$, $k_{cs}=0,4$.
  2. 2.
    Uncorrelated shift: the shifted spectrum is uncorrelated with the original spectrum. For $L=6$ and $N=8$, $k_{us}=1,2,3,5,6,7$.
Since (5) will be used to calculate the output PSD, $Y_{corr}(\Omega)$ needs to be pre-decorrelated before feeding into the system. By inspection, the sequence of correlated shift is
$k_{cs}[n]=\begin{cases}1,&\text{if mod}(n,N/\mathrm{gcd}(L,N))=0\\ {0},&{\text{else}}\end{cases}.$ (27)
Therefore the decorrelation sequence in (19) is chosen to be the “complement” of the correlated shift sequence (27), and the magnitude is set to preserve the power after decorrelation.